OpenStudy (anonymous):

can someone check this i got D please help me X = 18x+28 ---------------- 2 (x+14) Which of the following is a solution for x? A.) x=7 B.)x=-2 C.)x=-14 D.)x=2

OpenStudy (toxicsugar22):

hi can u help me

OpenStudy (anonymous):

D.

OpenStudy (toxicsugar22):

after her

OpenStudy (anonymous):

Thank you

OpenStudy (toxicsugar22):

hi can anyone help me after her

OpenStudy (anonymous):

\[x-\frac{18x+28}{2(x+14)}=0 \] Combine fractions into one. \[\frac{x^2+5 x-14}{x+14}=0 \] Numerator can be factored. \[\frac{(x-2) (x+7)}{x+14} \] x=2 and x=-7 will make the left hand side of the equation zero.